# Computation of Greeks under rough Volterra stochastic volatility models using the Malliavin calculus approach

### v1

@@ 2.3 Rough Volterra processes
Different kernels $K$ give different Gaussian Volterra processes. The most famous example is the standard fractional Brownian motion (fBm) $B^{H}_{t}$, that is a process with stationary increments and that corresponds to the representation
@@
where $K_{H}(t,s)$ is a quite complicated kernel that depends on the so-called Hurst parameter $H\in(0,1)$. See for example Nualart (2006).
@@
Other kernels $K_{H}(t,s)$ give the autocovariance function (15) despite they loose the property of stationary increments. In the present paper, for modeling purposes, and following many references in the literature, see for example Alòs, Mazet, and Nualart (2000)), we consider the so called simplified Riemann-Liouville kernel
$K_{H}(t,s)=\sqrt{2H}(t-s)^{H-1/2}.$ (16)
Recall that neither the standard fBm nor the simplified Riemann-Liouville process are semimartingales, see for example Thao (2006). This motivates the use of the so-called approximate fractional Brownian motion (afBm), i.e. a process with Volterra kernel
${K}_{H}(t,s)=\sqrt{2H}(t-s+\varepsilon)^{H-1/2},\quad\varepsilon\geq 0,\quad H\in(0,1).$ (17)
For every $\varepsilon>0$ such a process is a semimartingale and as $\varepsilon$ tends to zero, $B_{t}^{\varepsilon}$ converges to the Riemann-Liouville process $B_{t}$ in the $L^{2}-$norm, uniformly in $t\in[0,T]$ (Thao, 2006). In this case
@@
$\kappa_{t}:=\int_{0}^{t}{K}_{H}(t,s)\,\mathrm{d}s=\sqrt{2H}\int_{0}^{t}(t-s+\varepsilon)^{H-1/2}\,\mathrm{d}s$
@@
$\frac{\partial}{\partial H}{K}_{H}(t,s)={K}_{H}(t,s)\cdot\left(\frac{1}{2H}+\ln(t-s+\varepsilon)\right).$
@@
To avoid confusion, all theoretical calculations will be provided with a general kernel $K_{H}(t,s)$ and for numerical purposes the kernel (17) will be considered.
@@ 3.1 General SV model
where $r\geq 0$ is the constant instantaneous interest rate and $V$ denotes the volatility or the variance process that we assume adapted to the completed filtration generated by $W$ and $\widetilde{W}.$ Processes $S$ and $V$ are assumed to be square-integrable with continuous trajectories and such that $S_{t}$ and $V_{t}$ belong to ${\mathbb{D}}^{2,2}$ for any $t\in[0,T].$ Here ${\mathbb{D}}^{2,2}$ is the Sobolev space defined in the previous subsection associated to the Malliavin derivative with respect process $W$, the driving process of the price process. Note that $V_{t}$ can be seen as a functional of $W$ depending on an independent source of randomness $\widetilde{W}$ that can be treated as deterministic, see Nualart (2006).
@@
The following lemmas give the Malliavin derivatives of $S_{T}$ in terms of the Malliavin derivatives of $V_{t}.$ All the proofs are quite straightforward using that Itô integral in (19) can be seen as a Skorohod integral and using Proposition 2.3. Similar computations can be found for example in El-Khatib (2009) and Yolcu-Okur, Sayer, Yilmaz, and Inkaya (2018).
Similarly as in (Yolcu-Okur, Sayer, Yilmaz, and Inkaya, 2018, Def. 3, Prop. 6), we have the following result
@@
The proof is a straightforward computation of the Malliavin derivative of (19) using the Proposition 2.3. ∎
@@
Let $\theta$ now be a model parameter and our goal is to calculate the corresponding Greek, in particular the partial derivative with respect to it, see Definition 2.1. For example from (19) and using the Fubini’s theorem we may write
@@
Direct consequence of the IBP formula applied to the constant process $h_{t}\equiv\mathbf{1}_{[0,T]}(t)$ and Lemma 3.1 then gives us the general Greek formula
$\partial_{\theta}\mathbb{E}[\mathrm{e}^{-rT}f(S_{T}^{\theta})]$ $=\mathrm{e}^{-rT}\mathbb{E}\left[f(S_{T}^{\theta})\cdot\delta\left(\frac{\partial_{\theta}S_{T}^{\theta}}{\int_{0}^{T}S_{T}^{\theta}G(u,T)\,\mathrm{d}u}\right)\right]+\mathbb{E}\left[f(S_{T}^{\theta})\cdot\partial_{\theta}\left(\mathrm{e}^{-rT}\right)\right]$
$=\mathrm{e}^{-rT}\mathbb{E}\left[f(S_{T}^{\theta})\cdot\delta(c(V_{s}))\right]+\mathbb{E}\left[f(S_{T}^{\theta})\cdot\partial_{\theta}\left(\mathrm{e}^{-rT}\right)\right],$ (22)
@@
$c(V_{s})=\dfrac{\int_{0}^{T}b(V_{s})\,\mathrm{d}s}{\int_{0}^{T}G(u,T)\,\mathrm{d}u+\int_{0}^{T}a(V_{s})\,\mathrm{d}W_{s}}.$ (23)
@@
that depend on the model chosen for the volatility process is the essential problem in order to find Malliavin weights and to compute the considered Greeks. More concretely, the two objects depending on the model are $G(t,T)$ and $D_{s}G(t,T)$ with $s\leq t$.
@@
Let $G(t,T)$ be defined as in (20). Then the formulas for Greeks Delta, Gamma and Rho respectively are
$\Delta$ $=\mathbb{E}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{x}S_{T}}{\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u}\right)\right]$
$=\frac{e^{-rT}}{S_{0}}\mathbb{E}\left[f(S_{T})\left(\frac{W_{T}}{\int_{0}^{T}G(t,T)\,\mathrm{d}t}+\frac{\int_{0}^{T}\int_{s}^{T}D_{s}G(t,T)\,\mathrm{d}t\,\mathrm{d}s}{(\int_{0}^{T}G(t,T)\,\mathrm{d}t)^{2}}\right)\right],$
$\Gamma$ $=\mathbb{E}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{x}S_{T}}{\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u}\delta\left(\frac{\mathrm{e}^{-rT}\partial_{x}S_{T}}{\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u}\right)\right)\right]$
$\quad+\mathbb{E}\left[f(S_{T})\delta\left(\frac{\mathrm{e}^{-rT}\partial_{xx}S_{T}-\mathrm{e}^{-rT}\partial_{x}S_{T}\int_{0}^{T}D_{u}(\partial_{x}S_{T})\,\mathrm{d}u}{\left(\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u\right)^{2}}\right)\right]$
$=\frac{e^{-2rT}}{S_{0}^{2}}\mathbb{E}\left[\frac{f(S_{T})}{\int_{0}^{T}G(t,T)}\left(\frac{W_{T}}{\int_{0}^{T}G(t,T)\,\mathrm{d}t}+\frac{\int_{0}^{T}\int_{s}^{T}D_{s}G(t,T)\,\mathrm{d}t\,\mathrm{d}s}{(\int_{0}^{T}G(t,T)\,\mathrm{d}t)^{2}}\right)\right]$
$-\frac{e^{-rT}}{S_{0}^{2}}\mathbb{E}\left[f(S_{T})\left(\frac{W_{T}}{\int_{0}^{T}G(t,T)\,\mathrm{d}t}+\frac{\int_{0}^{T}\int_{s}^{T}D_{s}G(t,T)\,\mathrm{d}t\,\mathrm{d}s}{(\int_{0}^{T}G(t,T)\,\mathrm{d}t)^{2}}\right)\right]$
$\varrho$ $=\mathbb{E}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{r}S_{T}}{\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u}\right)\right]-\mathbb{E}\left[f(S_{T})\mathrm{e}^{-rT}T\right],$
[…]
@@
Follows directly from the Definition 2.1 and Theorem 2.4.
For the particular case of $\Delta,$ for example, we have $\partial_{S_{0}}S_{T}=\frac{S_{T}}{S_{0}}$ and then,
@@
$\partial_{S_{0}}\mathbb{E}[e^{-rT}f(S_{T})]=\frac{e^{-rT}}{S_{0}}\mathbb{E}\left[f(S_{T})\cdot\delta\left(\frac{1}{\int_{0}^{T}G(u,T)\,\mathrm{d}u}\right)\right].$ (26)
Applying formula (9) we obtain
@@
$\partial_{S_{0}}\mathbb{E}[e^{-rT}f(S_{T})]$ $=$ $\frac{e^{-rT}}{S_{0}}\mathbb{E}\left[f(S_{T})\cdot\delta\left(\frac{1}{\int_{0}^{T}G(u,T)\,\mathrm{d}u}\right)\right]$
@@
$=$ $\frac{e^{-rT}}{S_{0}}\mathbb{E}\left[f(S_{T})\left(\frac{W_{T}}{\int_{0}^{T}G(t,T)\,\mathrm{d}t}+\frac{\int_{0}^{T}\int_{t}^{T}D_{t}G(u,T)\,\mathrm{d}u\,\mathrm{d}t}{(\int_{0}^{T}G(t,T)\,\mathrm{d}t)^{2}}\right)\right].$
@@
Until now we did not have to specify a particular volatility process. From now on we will consider several particular cases of rough Volterra volatility processes. Having a particular volatility process will allow us to calculate also the Greeks $\mathcal{V}$ and $\mathcal{H}$ that were also introduced in Definition 2.1.

### v2

@@ A Gaussian Volterra process is defined as a process $Y=\{Y_{t},t\geq 0\}$ that can be represented as
Different kernels $K$ give different Gaussian Volterra processes. The most famous example is the standard fractional Brownian motion (fBm) that is a process with stationary increments and that corresponds to the representation
@@
where $K_{H}(t,s)$ is a quite complicated kernel that depends on the so-called Hurst parameter $H\in(0,1)$. See for example Nualart (2006) (Chapter 5).
@@
Other kernels $K_{H}(t,s)$ give the autocovariance function (21) despite they loose the property of stationary increments. In the present paper, for modeling purposes, and following many references in the literature, see for example Alòs, Mazet, and Nualart (2000), we consider the so called simplified Riemann-Liouville kernel
$K_{H}(t,s):=\sqrt{2H}(t-s)^{H-1/2}.$ (22)
Recall that neither the standard fBm nor the simplified Riemann-Liouville process are semimartingales, see for example Thao (2006). This motivates the use of the so-called approximate fractional Brownian motion (afBm), i.e. a Gaussian Volterra process with kernel
${K}_{H}(t,s):=\sqrt{2H}(t-s+\varepsilon)^{H-1/2},\quad\varepsilon\geq 0,\quad H\in(0,1).$ (23)
For every $\varepsilon>0$ such a process is a semimartingale and as $\varepsilon$ tends to zero it converges to the simplified Riemann-Liouville process in the $L^{2}-$norm, uniformly in $t\in[0,T];$ see Thao (2006). In this case
@@
$\kappa_{t}:=\int_{0}^{t}{K}_{H}(t,s)\,\mathrm{d}s=\sqrt{2H}\int_{0}^{t}(t-s+\varepsilon)^{H-1/2}\,\mathrm{d}s,$
$\frac{\partial}{\partial H}{K}_{H}(t,s)={K}_{H}(t,s)\cdot\left(\frac{1}{2H}+\ln(t-s+\varepsilon)\right),$
@@
$\kappa_{t}^{\prime}=\int_{0}^{t}\partial_{H}K_{H}(t,s)\,\mathrm{d}s.$
@@
To avoid confusion, all theoretical calculations will be provided with a general kernel $K_{H}(t,s)$ and for numerical purposes, the kernel (23) with a suitable value of $\varepsilon$ (typically $10^{-6}$) will be considered.
@@ 3.1 General SV model
where $r\geq 0$ is the constant instantaneous interest rate, $\sigma\in C^{2}(\mathbb{R})$ and $V$ denotes the volatility or the variance process that we assume adapted to the completed filtration generated by $W$ and $\widetilde{W}.$ Processes $S$ and $V$ are assumed to have continuous trajectories and such that $S_{t}$ and $V_{t}$ belong to ${\mathbb{D}}^{2,1}$ and ${\mathbb{D}}^{2,2}$ respectively for any $t\in[0,T].$ Here ${\mathbb{D}}^{2,p}$ is the Sobolev space defined in the previous section associated to the Malliavin derivative with respect process $W$, the driving process of the price process. Note that $V_{t}$ can be seen as a functional of $W$ depending on an independent source of randomness $\widetilde{W}$ that can be treated as deterministic, see Nualart (2006) (page 31).
@@
The following lemmas give the Malliavin derivatives of $S_{T}$ in terms of the Malliavin derivatives of $V_{t}.$ All the proofs are quite straightforward using that the Itô integral in (25) can be seen as a Skorohod integral and using Proposition 2.6. Similar computations can be found for example in El-Khatib (2009) and Yolcu-Okur, Sayer, Yilmaz, and Inkaya (2018).
Similarly as in Yolcu-Okur, Sayer, Yilmaz, and Inkaya (2018) (Def. 3, Prop. 6), we have the following result
@@
The proof is a straightforward computation of the Malliavin derivative of (25) using Proposition 2.6. ∎
@@
Let $\theta$ be a model parameter. Our goal is to calculate the corresponding Greek, in particular the partial derivative with respect to it, see Definition 2.1. For example from (25) and using Fubini’s theorem we may write
@@
Direct consequence of the IBP formula applied to a suitable process $h$ and Lemma 3.1 then gives us the general Greek formula
$\partial_{\theta}\mathbb{E}_{\mathbb{Q}}[\mathrm{e}^{-rT}f(S_{T}^{\theta})]$ $=\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T}^{\theta})\cdot\delta\left(\frac{\partial_{\theta}S_{T}^{\theta}\cdot h}{\int_{0}^{T}S_{T}^{\theta}G(u,T)h_{u}\,\mathrm{d}u}\right)\right]+\mathbb{E}_{\mathbb{Q}}\left[f(S_{T}^{\theta})\cdot\partial_{\theta}\left(\mathrm{e}^{-rT}\right)\right]$
$=\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T}^{\theta})\cdot\delta(c(V,h))\right]+\mathbb{E}_{\mathbb{Q}}\left[f(S_{T}^{\theta})\cdot\partial_{\theta}\left(\mathrm{e}^{-rT}\right)\right],$ (28)
@@
$c(V,h)=\dfrac{\int_{0}^{T}b(V_{s})\,\mathrm{d}s+\int_{0}^{T}a(V_{s})\,\mathrm{d}W_{s}}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u}\cdot h.$ (29)
@@
that depend on the model chosen for the volatility process is one of the the essential problems in order to find Malliavin weights and to compute the considered Greeks. More concretely, the two objects depending on the model are $G(t,T)$ and $D_{s}G(t,T)$ with $s\leq t$.
@@
Let $G(t,T)$ be defined as in (26). Then the formulas for Greeks Delta, Gamma, Rho, Vega and the derivative with respect to $H$ respectively are
$\Delta$ $=\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{S_{0}}S_{T}\cdot h}{\int_{0}^{T}D_{u}S_{T}h_{u}\,\mathrm{d}u}\right)\right]$
$=\frac{\mathrm{e}^{-rT}}{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{h}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u}\right)\right],$
$\Gamma$ $=\partial_{S_{0}}\frac{\mathrm{e}^{-rT}}{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{h}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u}\right)\right]$
$=-\frac{\mathrm{e}^{-rT}}{S_{0}^{2}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{h}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u}\right)\right]$
$+\frac{\mathrm{e}^{-rT}}{S_{0}^{2}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{h}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u}\delta\left(\frac{h}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u}\right)\right)\right]$
$\varrho$ $=\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{r}S_{T}\cdot h}{\int_{0}^{T}D_{u}S_{T}h_{u}\,\mathrm{d}u}\right)\right]-\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\mathrm{e}^{-rT}T\right],$
$=-T\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}[f(S_{T})]+T\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{h}{\int_{0}^{T}G(u,T)h_{u}\,\mathrm{d}u})\right)\right],$
$\mathcal{V}$ $=\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{\mathrm{e}^{-rT}\partial_{V_{0}}S_{T}\cdot h}{\int_{0}^{T}D_{u}S_{T}h_{u}\,\mathrm{d}u}\right)\right]=\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}[f(S_{T})\cdot\delta(c_{V_{0}}(V,h))],$
$\mathcal{H}$ $=\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\delta\left(\frac{\mathrm{e}^{-rT}\partial_{H}S_{T}\cdot h}{\int_{0}^{T}D_{u}S_{T}h_{u}\,\mathrm{d}u}\right)\right]=\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}[f(S_{T})\cdot\delta(c_{H}(V,h))].$
@@
We will show a formula for a general Greek of the form
$\partial_{\theta}\mathbb{E}_{\mathbb{Q}}[f(F^{\theta})\Phi^{\theta}].$
The conclusion of the theorem will follow by replacing $F^{\theta}=S_{T}$ and $\Phi^{\theta}=\mathrm{e}^{-rT}$.
Without loss of generality, assume that we are working under the canonical probability space, i.e. $\Omega=\mathcal{C}_{0}([0,T])$ is the space of continuous functions in $[0,T]$ vanishing at the origin endowed with the supremum norm $||\cdot||_{0}$, thus making $\Omega$ a Banach space. In other words, for $\omega\in\Omega$, $||\omega||_{0}=\sup_{t\in[0,T]}|\omega(t)|$. Let $\mathcal{F}$ the Borel $\sigma$-field of $\Omega$ and let $P$ be the probability measure such that the canonical process $W_{t}(\omega)=\omega(t)$ is a Brownian motion.
Given two sequences of positive real numbers $\{\delta_{n};n\geq 1\}$, $\{M_{n},n\geq 1\}$ such that $\delta_{n}\to 0$ in a monotonically decreasing way and $M_{n}\to\infty$ in a monotonically increasing way, we can define the sets $A^{j}_{n}$ for $j=1,2,3,4,5$ and $A_{n}^{\theta}$ as follows:
$A^{1.\theta}_{n}=\left\{\omega\in\Omega;\sup_{\omega^{\prime}\in\Omega;||\omega^{\prime}||_{0}\leq 1/n}\delta_{n}<F^{\theta}(\omega+\omega^{\prime})<M_{n}\right\},$
$A^{2,\theta}_{n}=\left\{\omega\in\Omega;\sup_{\omega^{\prime}\in\Omega;||\omega^{\prime}||_{0}\leq 1/n}\delta_{n}<|\partial_{\theta}F^{\theta}(\omega+\omega^{\prime})|<M_{n}\right\},$
$A^{3,\theta}_{n}=\left\{\omega\in\Omega;\sup_{\omega^{\prime}\in\Omega;||\omega^{\prime}||_{0}\leq 1/n}\delta_{n}<\left|\int_{0}^{T}D_{t}F^{\theta}(\omega+\omega^{\prime})\mathrm{d}t\right|<M_{n}\right\},$
@@
$A^{4,\theta}_{n}=\left\{\omega\in\Omega;\sup_{\omega^{\prime}\in\Omega;||\omega^{\prime}||_{0}\leq 1/n}\delta_{n}<\left|\int_{0}^{T}D_{t}\partial_{\theta}F^{\theta}(\omega+\omega^{\prime})\mathrm{d}t\right|<M_{n}\right\},$
$A^{5,\theta}_{n}=\left\{\omega\in\Omega;\sup_{\omega^{\prime}\in\Omega;||\omega^{\prime}||_{0}\leq 1/n}\delta_{n}<\left|D_{t}\int_{0}^{T}D_{s}F^{\theta}(\omega+\omega^{\prime})\mathrm{d}s\right|<M_{n}\text{ for all }t\in[0,T]\right\},$
$A^{\theta}_{n}=A^{1,\theta}_{n}\cap A^{2,\theta}_{n}\cap A^{3,\theta}_{n}\cap A^{4,\theta}_{n}\cap A^{5,\theta}_{n}.$
We also define the random variable
$\rho_{A^{\theta}_{n}}(\omega)=\inf\{||\omega^{\prime}||_{0};\omega+\omega^{\prime}\in A^{\theta}_{n}\}.$
Consider the smooth function $\phi(t)\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ satisfying that $\phi(t)=1$ in $|t|\leq\frac{1}{3}$, $\phi(t)=0$ if $|t|\geq\frac{2}{3}$ and $|\phi^{\prime}(t)|\leq 4$ for all $t\in\mathbb{R}$. Then, we define the sequence of random variables
$F_{n}^{\theta}=F^{\theta}\phi(n\rho_{A^{\theta}_{n}}).$
The first claim we want to show is that $\partial_{\theta}\phi(n\rho_{A^{\theta}_{n}}(\omega))=0$ almost surely. To see this, fix $\theta_{0}$ and assume for the sake of simplicity that there exists $\omega^{\prime}\in\Omega$ such that $||\omega^{\prime}||_{0}=\rho_{A_{n}^{\theta_{0}}}(\omega)$ (the general claim follows by replacing the existence of $h$ by the existence of a sequence of elements $\{\omega^{\prime}_{k};k\geq 1\}$ such that $||\omega^{\prime}_{k}||_{0}\to\rho_{A_{n}^{\theta_{0}}}(\omega)$ but the argument is identical). Now, since $||\omega^{\prime}||_{0}=\rho_{A_{n}^{\theta_{0}}}(\omega)$, it holds that $\omega+\omega^{\prime}\in A_{n}^{\theta_{0}}$. Now, due to the definition of the set $A_{n}^{\theta_{0}}$ and the continuity with respect to $\theta$ of all the random variables involved we conclude that $\omega+\omega^{\prime}\in A_{n}^{\theta}$ for all $\theta$ in a neighborhood of $\theta_{0}$. Since $\rho_{A_{n}^{\theta_{0}}}$ is defined as an infimum we deduce that
$\rho_{A_{n}^{\theta_{0}}}(\omega)\geq\rho_{A_{n}^{\theta}}(\omega)$
for all $\theta$ in a neighborhood of $\theta_{0}$. Since the choice of $\theta_{0}$ is arbitrary, we conclude that the function $\rho_{A_{n}^{\theta}}(\omega)$ is locally constant almost surely and therefore $\partial_{\theta}\rho_{A_{n}^{\theta}}(\omega)=0$ for every $\omega\in\Omega$.
Then, by the hypothesis on $F^{\theta}$ and following the same argument as in (Nualart, 2006, p. 231) we have that $A_{n}^{\theta}\uparrow\Omega$ as $n\to\infty$, $F_{n}^{\theta}=F^{\theta}$, $\partial_{\theta}F^{\theta}_{n}=\partial_{\theta}F^{\theta}$ on $A_{n}^{\theta}$ and $\partial_{\theta}F^{\theta}_{n}\in\mathbb{D}^{1,2}$, $F^{\theta}_{n}\in\mathbb{D}^{2,2}$. Consider now a localized Greek of the form
$\partial_{\theta}\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\Phi^{\theta}\phi(n\rho_{A^{\theta}_{n}})\right]=\mathbb{E}_{\mathbb{Q}}\left[f^{\prime}(F^{\theta})\partial_{\theta}F^{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})\right]+\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\partial_{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})\right].$
Applying the IBP formula in Theorem 2.7 to the first term we obtain that
@@
$\partial_{\theta}\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\Phi^{\theta}\phi(n\rho_{A^{\theta}_{n}})\right]=\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\delta\left(\frac{\partial_{\theta}F^{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\mathrm{d}t}\right)\right]+\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\partial_{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})\right].$
Notice that, by definition of $\phi(n\rho_{A_{n}^{\theta}})$, all the terms in the right-hand-side of the equation are well defined. We now will deduce the general formula for the Greeks using a limiting procedure. On the one hand, since $\phi(n\rho_{A_{n}^{\theta}})\to 1$ almost surely and it is bounded by $1$ we obtain that
$\lim_{n\to\infty}$ $\mathbb{E}_{\mathbb{Q}}\left[f^{\prime}(F^{\theta})\partial_{\theta}F^{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})\right]+\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\partial_{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})\right]$
$=$ $\mathbb{E}_{\mathbb{Q}}\left[f^{\prime}(F^{\theta})\partial_{\theta}F^{\theta}\Phi^{\theta}\right]+\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\partial_{\theta}\Phi^{\theta}\right]$
$=$ $\partial_{\theta}\mathbb{E}[f(F^{\theta})\Phi^{\theta}].$
Moreover, since $\phi(n\rho_{A_{n}^{\theta}})\to 1$, almost surely, the limit candidate (if it exists) of the sequence
@@
$\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\delta\left(\frac{\partial_{\theta}F^{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\mathrm{d}t}\right)\right]$
is
$\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\delta\left(\frac{\partial_{\theta}F^{\theta}\Phi^{\theta}h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\mathrm{d}t}\right)\right].$
Now, using that
$\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\delta\left(\frac{\partial_{\theta}F^{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\mathrm{d}t}\right)\right]=\mathbb{E}_{\mathbb{Q}}\left[f^{\prime}(F^{\theta})\partial_{\theta}F^{\theta}\Phi^{\theta}\phi(n\rho_{A_{n}^{\theta}})\right]$
and recalling the fact that the right-hand-side sequence is convergent we find that the left-hand-side sequence must be convergent. By uniqueness of the limit of sequences, we conclude that
$\partial_{\theta}\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\Phi^{\theta}\right]=\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\delta\left(\frac{\partial_{\theta}F^{\theta}\Phi^{\theta}h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\mathrm{d}t}\right)\right]+\mathbb{E}_{\mathbb{Q}}\left[f(F^{\theta})\partial_{\theta}\Phi^{\theta}\right]$
as desired. ∎
Remark 3.4.
In general, finding bounds for the Greeks which are independent of the model is not an easy problem. However, for the $\Delta$ case one can easily derive the estimate $0\leq\Delta\leq 1$ for Call options and $-1\leq\Delta\leq 0$ for Put options. Indeed, assume the payoff is $f(S_{T})=(S_{T}-K)^{+}$. Notice that we can write $S_{T}$ as
$S_{T}=S_{0}\exp\left(rT-\int_{0}^{T}\frac{\sigma(V_{s})^{2}}{2}\,\mathrm{d}s+\int_{0}^{T}\sigma(V_{s})\,\mathrm{d}W_{s}\right)=:S_{0}\mathcal{E}_{T}.$
Since $\mathcal{E}_{T}$ does not depend on $S_{0}$, clearly $\partial_{S_{0}}S_{T}=\mathcal{E}_{T}=\frac{S_{T}}{S_{0}}$ a.s. Now, by definition, the Delta can be computed as
$\Delta=\partial_{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[e^{-rT}(S_{T}-K)^{+}\right]$
where
$(S_{T}-K)^{+}=\begin{cases}S_{T}-K&S_{0}\geq\frac{K}{\mathcal{E}_{T}},\\ 0&S_{0}<\frac{K}{\mathcal{E}_{T}}.\end{cases}$
Hence,
$\partial_{S_{0}}(S_{T}-K)^{+}=\partial_{S_{0}}S_{T}\mathbf{1}_{\{S_{0}\geq\frac{K}{\mathcal{E}_{T}}\}}=\frac{S_{T}}{S_{0}}\mathbf{1}_{\{S_{0}\geq\frac{K}{\mathcal{E}_{T}}\}}=\frac{S_{T}}{S_{0}}\mathbf{1}_{\{S_{T}\geq K\}}.$
Plugging this into the definition of the Delta we have
$\Delta=\frac{1}{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[e^{-rT}S_{T}\mathbf{1}_{\{S_{T}\geq K\}}\right]$
Let’s proceed with the estimates. On the one hand, since $S_{T}$ and $S_{0}$ are positive, we clearly have
$\Delta=\frac{1}{S_{0}}\mathbb{E}\left[e^{-rT}S_{T}\mathbf{1}_{\{S_{T}\geq K\}}\right]\geq 0.$
On the other hand, since $\mathbf{1}_{\{S_{T}\geq K\}}\leq 1$, we have
$\Delta=\frac{1}{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[e^{-rT}S_{T}\mathbf{1}_{\{S_{T}\geq K\}}\right]\leq\frac{1}{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[e^{-rT}S_{T}\right]=\frac{S_{0}}{S_{0}}=1$
because discounted stock prices are martingales. As a conclusion,
$0\leq\Delta\leq 1$
for a Call option. Doing the same procedure for $(K-S_{T})^{+}$ we find that, for a Put option,
$-1\leq\Delta\leq 0.$
@@
Until now we did not have to specify a particular volatility process. From now on we will consider several particular cases of rough Volterra volatility processes.